OpenStudy (anonymous):

Leonard Townsend sold merchandise totaling $17,960 at a 3 1/2% rate of commission. How much commission did he earn?

OpenStudy (jdoe0001):

\(\bf {\color{brown}{ 3\frac{1}{2}}}\%\ of\ 17960 \implies \cfrac{{\color{brown}{ 3\frac{1}{2}}}}{100}\cdot 17960\implies {\color{brown}{ 3\frac{1}{2}}}\div 100\cdot 17960\)

OpenStudy (anonymous):

So the answer is

OpenStudy (isaidavila):

17920*7/200=??

OpenStudy (jdoe0001):

you tell us

OpenStudy (anonymous):

Not sure how to do it

OpenStudy (jdoe0001):

hmmm 3 1/2 is just 3 and half, or 3 and 0.5 or 3.5

OpenStudy (jdoe0001):

\(\bf {\color{brown}{ 3.5}}\%\ of\ 17960 \implies \cfrac{{\color{brown}{ 3.5}}}{100}\cdot 17960\implies {\color{brown}{ 3.5}}\div 100\cdot 17960\)

OpenStudy (anonymous):

So I multiply 17,960x3.5